OpenStudy (anonymous):

f f(x) = x + 9, find the value of x if f(x) = 15.

OpenStudy (lgbasallote):

change x to 15 so you'll get 15 + 9

Parth (parthkohli):

If you input a number, then you'll get the output with 9 added to the number. This is what the function says.

OpenStudy (anonymous):

as what i said in your earlier question.. you will just substitute the value of x which is 15 then just add it with 9..

OpenStudy (anonymous):

If f(x)=x + 9 and also f(x)=15 than you can say: x + 9 = 15 Now you only need to subtract 9 from both sides and you got your x. You can not say x=15 You would do this if the problem would be to find f(15)=15 + 9 = 21 But that is not the problem that you present here so this would be the wrong thing to do. Also you couldn´t solve for x in that case because there is no x left anymore.